The height at which the acceleration due to gravity becomes g/9 in terms of R, the radius of earth is?
1.2R
2.R/root2
3.R/2
4.root2/R

Grade:9

1 Answers

ManpreetK
6 years ago
a =g/9 Let a=g`g`= g/(1+h/R) ^2g/9 = g/(1+h/R)^21/9 = 1/(1+h/R)^29= (1+h/R)^23= (1+h/R)h/R = 2h=2R